OpenStudy (anonymous):

what is 65% of 80 I did it two ways. Which is correct? 0.65 * x = 80 x = 80/0.65 x = 123.7 or 80 * .65 = 52

OpenStudy (anonymous):

80-100% x-65% x=80*65/100=52

OpenStudy (anonymous):

use proportion, but it seems second is correct

OpenStudy (anonymous):

When doing questions like this its often to use a common sense check. If you take 65% of something, the result will always be smaller than what you started with, so for example, your first answer cannot be right. Second answer is correct.

OpenStudy (anonymous):

Thought so; thanks bunches. Always good to ask. :)